# Large Language Model Psychometrics: A Systematic Review of Evaluation, Validation, and Enhancement

### v1

@@ 5.1.2 Values
"Values are enduring beliefs that guide behavior and decision-making, reflecting what is important and desirable to an individual or group." [Schwartz, 1992]
Values are central to understanding human motivation, attitudes, and behavior, shaping how individuals perceive themselves and the world around them. According to Schwartz’s theory of basic human values [Schwartz, 2012], the most established value theory, they 1) are beliefs linked inextricably to affect, 2) are desirable goals that motivate actions, 3) transcend specific situations and actions, 4) serve as standards or criteria for evaluating and selecting actions, policies, people, and events, and 5) are ordered by relative importance, which guides choices and actions. These unique characteristics make values a powerful lens for understanding LLM behavior. For example, Ye et al. [2025a] show how different values contribute to the safety of LLMs; Liu et al. [2024b] reveal that different spiritual values affect LLMs in social-fairness scenarios; and Sorensen et al. [2024b] demonstrate that standard value alignment reduces distributional pluralism in LLM outputs.
Schwartz’s Theory.
Schwartz [1992] identify 10 basic human values: Self-Direction, Stimulation, Hedonism, Achievement, Power, Security, Conformity, Tradition, Benevolence, and Universalism. These values are further grouped into four higher-order dimensions: Openness to Change, Conservation, Self-Enhancement, and Self-Transcendence. Later iterations of the theory refine its nuances by subdividing the 10 values into finer categories [Schwartz et al., 2012]. Most research on Schwartz’s values in LLMs focuses on the 10 basic values, while some studies explore the 19 refined values [Rozen et al., 2024]. Many investigations directly apply established inventories, i.e., Schwartz Value Survey (SVS) [Schwartz, 1992] or the Portrait Values Questionnaire (PVQ) [Schwartz et al., 2001] (e.g., [Hadar-Shoval et al., 2024, Fischer et al., 2023, Kovač et al., 2023, Lee et al., 2024a, Miotto et al., 2022]). Other research adapts these inventories to better suit the context of LLMs. For example, Ren et al. [2024] rephrase survey items into advice-seeking questions, and Shen et al. [2024] transform subjective statements into opinion-seeking questions about AI actions. Rather than directly administering or rephrasing static inventories, Ye et al. [2025a], Chiu et al. [2025], Yao et al. [2024] leverage existing cross-domain LLM prompt datasets or LLM-generated contextualized questions as test items.
World Values Survey (WVS).
The World Values Survey (WVS) is a global research project that explores people’s values and beliefs, how they change over time, and their social and political impact [Haerpfer et al., 2022]. It measures a broad spectrum of values, including cultural values defined on two broad dimensions: traditional vs. secular-rational values and survival vs. self-expression values. Kim and Baek [2024] directly prompt LLMs with WVS survey items. In contrast, Yao et al. [2025a], Chiu et al. [2025] use model-generated or self-designed prompts to measure these cultural values in LLMs.
Values Survey Module (VSM).
Hofstede [1984] identifies six key value dimensions that describe cultural differences: Power Distance, Individualism vs. Collectivism, Masculinity vs. Femininity, Uncertainty Avoidance, Long-term vs. Short-term Orientation, and Indulgence vs. Restraint. These dimensions help explain how cultural values influence behavior in societies and organizations [Hofstede, 2001]. VSM is a questionnaire-based tool that assesses cultural preferences across these dimensions [Hofstede, 2011]. Kovač et al. [2023], Zhong et al. [2024] directly apply the VSM to evaluate cultural values in LLMs, while Ren et al. [2024], Kharchenko et al. [2024], Ye et al. [2025a] adapt the VSM or use self-designed inventories to measure these dimensions.
Global Leadership and Organizational Behavior Effectiveness (GLOBE).
The GLOBE study investigates cultural values, leadership behaviors, and organizational practices across countries [House et al., 2004]. Building on Hofstede’s work [Hofstede, 1984], it incorporates additional dimensions: Performance Orientation, Assertiveness, and Humane Orientation, resulting in nine cultural dimensions. Karinshak et al. [2024], Li et al. [2024d], Ren et al. [2024] apply or adapt the GLOBE culture questionnaire to evaluate cultural values in LLMs.
Social Value Orientation (SVO).
SVO is a psychological framework that measures individual preferences for allocating resources between oneself and others [Messick and McClintock, 1968]. It focuses on distinguishing between prosocial (e.g., altruists, cooperators) and proself (e.g., individualists, competitors) orientations. SVO is typically measured using experimental tasks or questionnaires, such as the SVO Slider Measure [Murphy et al., 2011] or Decomposed Games [Liebrand, 1984]. These tools assess how individuals allocate resources in hypothetical scenarios involving themselves and others. For example, Zhang et al. [2024c] apply the SVO Slider Measure to evaluate social preferences in LLMs.
Other or Custom Theories and Inventories.
Meadows et al. [2024], Xu et al. [2023] introduce localized inventories to evaluate LLMs based on Australian and Chinese values, respectively. Some researchers focus on specific topics, such as age-related values [Zhang et al., 2024b] and spiritual values [Liu et al., 2024b]. Other studies involve large-scale, cross-domain prompt datasets [Li et al., 2024d, Moore et al., 2024], among which Jiang et al. [2024a] concentrate on eliciting harmful, biased, or ethics-related content from LLMs. Notably, Biedma et al. [2024], Ye et al. [2025b] challenge established value theories and propose novel ones for LLM values, built in a bottom-up manner using psycholexical data.
Comparing Value Theories and Suggesting Use Cases.
@@
Cross-cultural research suggests that LLMs may embody a blend of cultural values, integrating perspectives from diverse backgrounds [Kovač et al., 2023]. However, they generally exhibit a tendency toward Western liberal values [Kim and Baek, 2024]. LLMs can display different values based on profiling prompts [Kharchenko et al., 2024, Zhong et al., 2024, Karinshak et al., 2024]. The context-dependency of their values is concerning and challenges the stability assumptions in human value theories [Shen et al., 2024, Meadows et al., 2024, Moore et al., 2024, Chiu et al., 2025, Xu et al., 2023, Kovač et al., 2023]. Further discussions in reliability and validity are presented in § 7.
@@ 5.1.3 Morality
"Morality is the categorization of intentions, decisions and actions into those that are proper, or right, and those that are improper, or wrong." [Long and Sedley, 1987]
Morality is a fundamental aspect of human behavior, influencing social interactions, decision-making, and ethical reasoning. The Moral Foundations Theory (MFT) [Graham et al., 2009] posits that morality is shaped by six innate psychological systems: Care/Harm, Fairness/Cheating, Loyalty/Betrayal, Authority/Subversion, Sanctity/Degradation, and Liberty/Oppression. These moral foundations are hypothesized to be universal across cultures, providing a common framework for understanding moral judgments and behavior. The theory suggests that individuals vary in the extent to which they endorse each foundation, leading to diverse moral profiles that influence their attitudes, beliefs, and actions.
It is crucial to conduct moral assessments of LLMs to ensure their ethical deployment. A large body of research (e.g., [Tlaie, 2024, Liu et al., 2024d, Fraser et al., 2022, Yao et al., 2025a, Zhou et al., 2024a]) applies MFT, primarily through the Moral Foundations Vignettes (MFVs) [Clifford et al., 2015], the Moral Foundations Questionnaire (MFQ) [Graham et al., 2009], the MFQ-2 [Atari et al., 2023], and the Moral Foundations Dictionary (MFD) [Graham et al., 2009]. These papers investigate model bias, alignment with political/moral ideologies, and variation in personal or cultural values, using controlled prompt tests [Tlaie, 2024, Nunes et al., 2024, Abdulhai et al., 2024], persona-driven exploration [Münker, 2024], or evaluating internal moral coherence [Nunes et al., 2024].
Other prominent moral theoretical frameworks and survey tools for evaluating LLMs include Kohlberg’s Theory via the Defining Issues Test (DIT) [Kohlberg, 1964], the Consequentialist-Deontological distinction [Beauchamp, 2001], and the PEW 2013 Global Attitudes Survey [Center, 2013]. Khandelwal et al. [2024], Tanmay et al. [2023] utilize the DIT, a psychometric tool aligned with Kohlberg’s stages of moral development, to analyze structural reasoning in moral dilemmas. They report that models like GPT-4 exhibit post-conventional moral reasoning. Khandelwal et al. [2024] further extend this analysis to a multilingual context, examining cross-linguistic variations in moral reasoning abilities. Neuman et al. [2025] explore ethical reasoning using established typologies, including consequentialist-deontological analysis. The results uncover the general rationalist, consequentialist emphasis in LLMs. According to the PEW 2013 Global Attitudes Survey, Meijer et al. [2024] find that the models display homogeneous moral values and have limited effectiveness in reflecting cross-cultural morality.
Localized moral theoretical frameworks are essential for ensuring the ethical deployment of LLMs in various cultural contexts. Liu et al. [2024d] develop the Chinese Moral Dictionary (CMD) and find that ChatGPT and Gemini prefer individualistic moral beliefs, while Ernie and ChatGLM lean towards collectivist moral beliefs. Takeshita et al. [2023], Ohashi et al. [2024] create the JCommonsenseMorality (JCM) dataset and fine-tune LLMs to adapt to Japanese culture.
Researchers have also developed specialized datasets for studying LLM morality. Hendrycks et al. [2021] create the ETHICS dataset, which is then utilized and extended to evaluate LLMs’ knowledge of morality, covering concepts such as justice, well-being, duties, virtues, and commonsense morality [Jinnai, 2024, Rodionov et al., 2023, Karpov et al., 2024, Yu et al., 2023, Albrecht et al., ]. Marraffini et al. [2024] develop the Greatest Good Benchmark (GGB) to evaluate LLMs’ moral judgments using utilitarian dilemmas. Jin et al. [2024b] develop the MultiTP, a cross-lingual corpus of moral dilemma vignettes in over 100 languages.
@@
5.1.4 Attitudes & Opinions
"Attitudes are always attitudes about something. This implies three necessary elements: first, there is the object of thought, which is both constructed and evaluated. Second, there are acts of construction and evaluation. Third, there is the agent, who is doing the constructing and evaluating. We can therefore suggest that, at its most general, an attitude is the cognitive construction and affective evaluation of an attitude object by an agent." [Bergman, 1998]
We use the term "attitude" to encompass both attitudes and opinions, following [Bergman, 1998, Ma et al., 2024a], who treat these concepts as synonymous. Most research on LLM attitudes examines political attitudes and public opinions [Ma et al., 2024a], which are key cognitive and behavioral foundations in human society and closely tied to model fairness, credibility, and social impact [Durmus et al., 2023, Santurkar et al., 2023, Sanders et al., 2023, Lee et al., 2024c, Hartmann et al., 2023]. Other biases—such as those related to gender, race, culture, religion, occupation and age—are also studied as part of LLM attitudes, though many works approach them from data and algorithmic perspectives. For a broader discussion on bias and fairness in LLMs, we refer readers to [Gallegos et al., 2024, Ranjan et al., 2024].
To measure the political attitudes of LLMs, researchers often employ standardized questionnaires and scales from political science and social psychology. A widely used tool is the American National Election Studies (ANES) [^1] https://electionstudies.org/about-us/., applied in works such as Qi et al. [2024], Jiang et al. [2022, 2024d], Yang et al. [2024] to assess LLMs’ stances on U.S. political issues. Similarly, the American Trends Panel (ATP) [^2] https://www.pewresearch.org/american-trends-panel/. has been leveraged by Santurkar et al. [2023], Hwang et al. [2023], Tjuatja et al. [2024] to analyze opinion distributions on public policy. For cross-national comparisons, studies like Ma et al. [2024b], Ball et al. [2025], von der Heyde et al. [2024] utilize the German Longitudinal Election Study (GLES) [^3] https://www.gesis.org/en/gles.. Additionally, the Political Compass Test (PCT) [^4] https://www.politicalcompass.org/test. has gained prominence for positioning LLMs within multidimensional political spectrums [Azzopardi and Moshfeghi, 2024, Bernardelle et al., 2024, Hartmann et al., 2023, Röttger et al., 2024].
Other survey instruments include the General Social Survey (GSS) [Kim and Lee, 2023], American Community Survey (ACS) [Dominguez-Olmedo et al., 2024], Canadian Election Study (CES) [Sanders et al., 2023], European Social Survey (ESS) [Geng et al., 2024], Survey of Russian Elites [Kalinin, 2023], and Supreme Court Case Political Evaluation (SCOPE) [Xu et al., 2025c]. Additionally, researchers have developed specialized datasets for LLMs, such as OpinionQA [Santurkar et al., 2023], a comprehensive dataset based on ATP surveys; IssueBench [Röttger et al., 2025], a benchmark covering controversial issues with multiple response formats; and GlobalOpinionQA [Durmus et al., 2023], which extends political opinion evaluation to cross-cultural contexts.
Researchers also present tailored tools for LLMs, due to the complexity of measuring political attitudes. Haller et al. [2023] introduce OpinionGPT, a web tool demonstrating how input data biases influence model outputs. Chalkidis and Brandl [2024] fine-tune LLMs on European Parliament speeches to assess political knowledge and reasoning, realigning the models to reflect specific stances. Coppolillo et al. [2025] propose a framework for quantifying biases in multi-agent generative systems by simulating echo chambers—paired LLMs with aligned perspectives discuss polarizing topics. Finally, Kim et al. [2025b] investigate linear interpolation methods to track continuous shifts in model political attitudes.
@@ 5.1.4 Attitudes & Opinions
Social Intelligence Leng and Yuan [2023], Mittelstädt et al. [2024], Shapira et al. [2023b], Zhou et al. [2024b], Wang et al. [2024d], Liu et al. [2024c], Wang et al. [2024a], Mou et al. [2024], Liang et al. [2024], Zadeh et al. [2019], Xu et al. [2024b], Guo et al. [2023a], Kovač et al. [2024, 2021], Mathur et al. [2024]
@@ 5.2.1 Heuristics and Biases
"Heuristics and biases are mental shortcuts or rules of thumb that simplify decision-making and problem-solving." [Tversky and Kahneman, 1974]
Heuristics shape both the strengths and biases in human decision-making. Recent research employs psychometric tools to systematically evaluate rationality and biases in LLM outputs and offer theoretical explanations for these biases.
Pioneering work by Binz and Schulz [2023] applies a battery of canonical cognitive ability tests to GPT-3. The test results demonstrate that the model performs on par with or better than human participants, while also exhibiting human-like cognitive biases such as framing effects, certainty effects, and overweighting biases. Subsequent studies include more advanced LLMs and reveal a dual-process dynamic [Hagendorff et al., 2023]: earlier GPT models exhibit more human-like System 1 errors, while more advanced models such as GPT-4 display more hyperrational System 2 reasoning and thus fewer biases. In a similar vein, but with an expanded scope, Yax et al. [2024], Sundaram and Alwar [2024], Momennejad et al. [2023], Echterhoff et al. [2024], Saeedi et al. [2024], Ranaldi and Zanzotto [2024], Macmillan-Scott and Musolesi [2024], Ando et al. [2023] evaluate LLMs regarding biases including the conjunction fallacy, unwarranted beliefs, anchoring bias, status-quo bias, loss aversion, positional selection biases, and Clever Hans effects. Later work significantly scales up the evaluation with synthetic items. Malberg et al. [2024] evaluate cognitive biases in 20 LLMs based on novel decision-making scenarios and a dataset of 30,000 tests covering 30 cognitive biases. Xie et al. [2024b] introduce the "MindScope" dataset and multi-agent methods to evaluate 72 bias categories using psychometric-inspired multi-round dialogue experiments. Recently, the evaluation paradigm has also been extended to multimodal models [Schulze Buschoff et al., 2025].
To understand the origins of these biases, researchers have proposed diverse theoretical frameworks. Hagendorff et al. [2023], Yax et al. [2024] link dual-process theory (System 1 vs. System 2 thinking) with the contextualized inference of LLMs, contrasting intuitive and deliberative reasoning modes. Scholten et al. [2024] explain LLM biases using metacognitive myopia, a syndrome described as being "relatively accurate in utilizing even large amounts of stimulus information, but naïve and almost blind regarding the history and validity of the stimulus data." The authors argue that LLMs lack metacognitive abilities such as monitoring and control, which result in five symptoms of metacognitive myopia and, therefore, systematic biases. Sundaram and Alwar [2024] explore cognitive dissonance theory and elaboration likelihood theory to explain inconsistencies in LLM reasoning. They suggest that certain biases in LLMs may not just be drawbacks but could serve as a means of identifying and mitigating logical fallacies in human reasoning. Differing from the cognitive psychology perspective, Shaikh et al. [2024] propose a mechanistic interpretability framework for sourcing the biases. They leverage influence graphs and Shapley value analysis to interpret biases and explore connections between model training and bias emergence.
@@ 5.2.2 Social Interactions
"Theory of Mind is the ability to attribute mental states such as beliefs, intentions, and knowledge to others." [Premack and Woodruff, 1978]
@@
Advanced LLMs can simulate ToM-like reasoning under certain conditions, which prompts questions about how these behaviors arise and how robustly they generalize. Foundational studies apply classic psychometrics—false belief tasks—to evaluate ToM in LLMs. [Kosinski, 2023a, b] report that GPT-3.5 and GPT-4 perform at levels matching or exceeding those of children in structured ToM tasks, suggesting the spontaneous emergence of ToM-like behavior as models scale in size. Through mechanistic interpretation, Jamali et al. [2023] offer evidence of parallels between the LLM embeddings and neurons in the human brain.
However, these claims have been challenged. Ullman [2023], Shapira et al. [2023a], Holterman and Deemter [2023] soon demonstrate that minor task perturbations often lead to poor performance, indicating reliance on brittle heuristics rather than genuine semantic comprehension. Hoping to address the controversy, researchers have developed robust benchmarks and evaluation protocols. For example, BigToM [Gandhi et al., 2023], ToMBench [Chen et al., 2024d], HI-TOM [He et al., 2023], OpenToM [Xu et al., 2024a], and ToMATO [Shinoda et al., 2025] feature procedurally generated, higher-order, or broader ToM tasks. These benchmarks often reveal performance drop-offs in more complex tasks, such as 6th-order belief attribution [Street et al., 2024] or detecting faux pas [Strachan et al., 2024a]. But others argue the failures result from a lack of more general commonsense reasoning, rather than a failure to represent mental states [Pi et al., 2024]. Riemer et al. [2025] also argue that we should distinguish between literal and functional ToM, and that current ToM benchmarks are inadequate for assessing functional ToM. Recent attempts also extend ToM evaluation to multimodal settings [Chen et al., 2024c, Strachan et al., 2024b, Jin et al., 2024a] and multi-agent interactions [Li et al., 2023b].
Several studies explore enhancement techniques. Prompt engineering methods like reasoning, reflection, and planning [Moghaddam and Honey, 2023, Lin et al., 2024, Zhou et al., 2023a, Yang et al., 2025a, Zhang et al., 2025]; persona-based prompts [Tan et al., 2024]; and psychology-informed prompts [Leer et al., 2023, Wilf et al., 2023] can significantly increase accuracy but raise questions about the authenticity of improvements. Others implement symbolic reasoning frameworks, such as SymbolicToM [Sclar et al., 2023, Xu et al., 2025a], ToM-LM [Tang and Belle, 2024], or task decomposition strategies [Sarangi et al., 2025], to enhance interpretability and belief state modeling. While recent models keep progressing in ToM tasks, especially in structured tasks [Strachan et al., 2024a, Street et al., 2024], they are shown to remain inconsistent in open-ended, adversarial, or pragmatic reasoning settings [Sclar et al., 2024, Yu et al., 2025, Nickel et al., 2024]. Comparisons with human baselines show that LLMs can approximate ToM behavior in narrow contexts but fall short of general human-like social cognition [Strachan et al., 2024a, van Duijn et al., 2023, Jones et al., 2024].
Main Findings.
@@
"Emotional Intelligence is the subset of social intelligence that involves the ability to monitor one’s own and others’ feelings and emotions, to discriminate among them and to use this information to guide one’s thinking and actions." [Salovey and Mayer, 1990]
@@
Recent studies introduce benchmark tools specifically designed to measure EI in LLMs through structured tasks informed by psychological theories. For example, EmoBench [Sabour et al., 2024] and EQ-Bench [Paech, 2023] provide task sets operationalizing constructs like emotional understanding and regulation, grounding their evaluations in frameworks such as the Mayer-Salovey-Caruso Emotional Intelligence Test (MSCEIT). Others adapt existing psychometric instruments to LLM psychometrics, including the Situational Evaluation of Complex Emotional Understanding (SECEU) [Wang et al., 2023b], the Levels of Emotional Awareness Scale (LEAS) [Elyoseph et al., 2023], the Toronto Alexithymia Scale (TAS-20) and Empathy Quotient (EQ-60) [Patel and Fan, 2023], and others [Huang et al., 2024b]. Models such as GPT-4 are shown to match or exceed human baseline in measures of emotional awareness and understanding [Wang et al., 2023b, Elyoseph et al., 2023, Patel and Fan, 2023], though it has been reported that LLMs lack deep reflexive analysis of emotional experiences and motivational aspects of emotions [Vzorinab et al., 2024]. Recent benchmark tools also extend EI evaluations to multimodal settings [Hu et al., 2025].
A subset of the literature further explores methodological pathways for improving LLM emotional competencies. Modular Emotional Intelligence (MoEI) [Zhao et al., 2024] and Emotional Chain-of-Thought (ECoT) [Li et al., 2024e] offer strategies to enhance emotionally intelligent behaviors while preserving general cognitive functionality. Another line of research explores LLMs’ ability to simulate nuanced emotional alignment, empathy, and scenario-appropriate reactions, based on Emotion Appraisal Theory [Roseman and Smith, 2001] and Emotion-Focused Theory [Greenberg, 2004]. Surprisingly, Li et al. [2023a] indicate that emotional stimuli improve general performance of LLMs.
Main Findings.
@@
"Social Intelligence is the ability to understand and manage people." [Thorndike and Stein, 1937]

### v2

@@ 4.1.2 Values
Values are enduring beliefs that guide behavior and decision-making, reflecting what is important and desirable to an individual or group [Schwartz, 1992]. Values offer a powerful lens for understanding LLM behavior. For example, Ye et al. [2025a] show how different values contribute to the safety of LLMs; Liu et al. [2024b] reveal that different spiritual values affect LLMs in social-fairness scenarios; and Sorensen et al. [2024b] demonstrate that standard value alignment reduces distributional pluralism in LLM outputs.
A growing body of research has applied diverse value theories and instruments to assess the value orientations of LLMs. Schwartz’s Value Theory, with its ten basic values and higher-order dimensions, is the most widely used, often via the Schwartz Value Survey (SVS) or Portrait Values Questionnaire (PVQ) [Schwartz, 1992, Schwartz et al., 2001]. Adaptations and contextualized prompts are increasingly common to better suit LLMs’ operational context [Ren et al., 2024, Shen et al., 2024]. The World Values Survey (WVS) has also been used to probe LLMs’ alignment with societal-level value dimensions [Haerpfer et al., 2022, Kim and Baek, 2024]. Cross-cultural frameworks like Hofstede’s Values Survey Module (VSM) and the GLOBE study extend this analysis to workplace and leadership-related cultural dimensions, with both direct and adapted inventories employed [Hofstede, 1984, House et al., 2004, Ren et al., 2024]. Social Value Orientation (SVO) frameworks, using tools like the SVO Slider Measure, focus on LLMs’ prosocial versus proself tendencies [Murphy et al., 2011, Zhang et al., 2024c]. Additionally, researchers have developed localized or custom inventories to capture region-specific or topical values, and some propose novel, bottom-up value taxonomies for LLMs using psycholexical data [Meadows et al., 2024, Xu et al., 2023, Biedma et al., 2024, Ye et al., 2025b]. Collectively, these studies reveal both the adaptability of human value frameworks to LLMs and the need for tailored approaches.
@@
Cross-cultural research suggests that LLMs may embody a blend of cultural values, integrating perspectives from diverse backgrounds [Kovač et al., 2023]. However, they generally exhibit a tendency toward Western liberal values [Kim and Baek, 2024]. LLMs can display different values based on profiling prompts [Kharchenko et al., 2024, Zhong et al., 2024, Karinshak et al., 2024]. The context-dependency of their values is concerning and challenges the stability assumptions in human value theories [Shen et al., 2024, Meadows et al., 2024, Moore et al., 2024, Chiu et al., 2025, Xu et al., 2023, Kovač et al., 2023].
Comparing value theories and suggesting use cases.
@@ 4.1.3 Morality
Morality is the categorization of intentions, decisions and actions into those that are proper, or right, and those that are improper, or wrong [Long and Sedley, 1987]. It is crucial to conduct moral assessments of LLMs to ensure their ethical deployment. A large body of research applies the Moral Foundations Theory (MFT) [Graham et al., 2009], primarily through the Moral Foundations Vignettes (MFVs) [Clifford et al., 2015], the Moral Foundations Questionnaire (MFQ) [Graham et al., 2009], the MFQ-2 [Atari et al., 2023], and the Moral Foundations Dictionary (MFD) [Graham et al., 2009]. These papers investigate model bias, alignment with political/moral ideologies, and variation in personal or cultural values, using controlled prompt tests [Tlaie, 2024, Nunes et al., 2024, Abdulhai et al., 2024], persona-driven exploration [Münker, 2024], or evaluating internal moral coherence [Nunes et al., 2024]. Other prominent moral theories and instruments include Kohlberg’s Theory via the Defining Issues Test (DIT) [Kohlberg, 1964], the Consequentialist-Deontological distinction [Beauchamp, 2001], the PEW 2013 Global Attitudes Survey [Center, 2013], and localized moral theories [Liu et al., 2024d, Takeshita et al., 2023]. Specialized datasets are also curated for scalable and comprehensive morality evaluation [Hendrycks et al., 2021, Jinnai, 2024, Marraffini et al., 2024, Jin et al., 2024b]
Main findings.
@@
4.1.4 Attitudes and opinions
Attitudes are always attitudes about something. This implies three necessary elements: first, there is the object of thought, which is both constructed and evaluated. Second, there are acts of construction and evaluation. Third, there is the agent, who is doing the constructing and evaluating. We can therefore suggest that, at its most general, an attitude is the cognitive construction and affective evaluation of an attitude object by an agent [Bergman, 1998]. We use the term "attitude" to encompass both attitudes and opinions, following [Bergman, 1998, Ma et al., 2024a]. Most research on LLM attitudes examines political attitudes and public opinions [Ma et al., 2024a], which are key cognitive and behavioral foundations in human society and closely tied to model fairness, credibility, and social impact [Durmus et al., 2023, Santurkar et al., 2023, Sanders et al., 2023, Lee et al., 2024c, Hartmann et al., 2023]. For a broader discussion on bias and fairness in LLMs, we refer readers to [Gallegos et al., 2024, Ranjan et al., 2024].
Researchers assess LLMs’ political attitudes using standardized questionnaires and scales from political science and social psychology. US-focused instruments include the American National Election Studies (ANES) [Qi et al., 2024] and the American Trends Panel (ATP) [Santurkar et al., 2023]. For cross-national comparisons, studies use tools like the German Longitudinal Election Study (GLES) [Ma et al., 2024b, Ball et al., 2025, von der Heyde et al., 2024]. The Political Compass Test (PCT) is also widely used to map LLMs within multidimensional political spectrums [Azzopardi and Moshfeghi, 2024, Bernardelle et al., 2024, Hartmann et al., 2023, Röttger et al., 2024].
Other survey instruments include the General Social Survey (GSS) [Kim and Lee, 2023], the American Community Survey (ACS) [Dominguez-Olmedo et al., 2024], the Canadian Election Study (CES) [Sanders et al., 2023], the European Social Survey (ESS) [Geng et al., 2024], the Survey of Russian Elites [Kalinin, 2023], and the Supreme Court Case Political Evaluation (SCOPE) [Xu et al., 2025c]. Additionally, researchers have developed specialized datasets and tailored tools for LLMs, such as OpinionQA [Santurkar et al., 2023], a comprehensive dataset based on ATP surveys; IssueBench [Röttger et al., 2025], a benchmark covering controversial issues with multiple response formats; GlobalOpinionQA [Durmus et al., 2023], which extends political opinion evaluation to cross-cultural contexts; and OpinionGPT [Haller et al., 2023], a web tool that demonstrates how input data biases influence model outputs.
@@ 4.2 Measuring cognitive constructs
Social Intelligence Leng and Yuan [2023], Mittelstädt et al. [2024], Shapira et al. [2023b], Zhou et al. [2024b], Wang et al. [2024d], Liu et al. [2024c], Wang et al. [2024a], Mou et al. [2024], Liang et al. [2024], Zadeh et al. [2019], Xu et al. [2024b], Guo et al. [2023a], Kovač et al. [2024, 2021], Mathur et al. [2024], Schoenegger et al. [2025]
@@ 4.2.1 Heuristics and biases
Heuristics and biases are mental shortcuts that simplify decision-making but can introduce systematic errors [Tversky and Kahneman, 1974]. Recent research employs psychometric tools to evaluate rationality and biases in LLM outputs and offer theoretical explanations for these biases.
Early studies, such as Binz and Schulz [2023], found that GPT-3 performs comparably to humans on cognitive ability tests, exhibiting biases like framing and anchoring. More advanced models, such as GPT-4, show fewer System 1 (intuitive) errors and more System 2 (deliberative) reasoning, though biases persist [Hagendorff et al., 2023]. Large-scale evaluations now cover a wide range of biases—including conjunction fallacy, unwarranted beliefs, and loss aversion—using synthetic scenarios and multi-agent dialogue [Malberg et al., 2024, Xie et al., 2024b]. Multimodal models are also being assessed for similar patterns [Schulze Buschoff et al., 2025].
To explain these findings, researchers draw on dual-process theory, linking LLM reasoning to intuitive and deliberative modes [Hagendorff et al., 2023, Yax et al., 2024]. Other frameworks highlight LLMs’ lack of metacognitive abilities, leading to “metacognitive myopia”—the inability to monitor or control reasoning processes, which perpetuates systematic biases [Scholten et al., 2024]. Additional perspectives use cognitive dissonance and elaboration likelihood theories to interpret inconsistencies [Sundaram and Alwar, 2024], while mechanistic approaches leverage influence graphs and Shapley values to trace the origins of biases in model training [Shaikh et al., 2024].
@@ 4.2.2 Social interactions
ToM is the ability to attribute mental states such as beliefs, intentions, and knowledge to others [Premack and Woodruff, 1978]. Advanced LLMs can simulate ToM-like reasoning under certain conditions, which prompts questions about how these behaviors arise and how robustly they generalize. Foundational studies apply classic psychometrics—false belief tasks—to evaluate ToM in LLMs. [Kosinski, 2023a, b] report that GPT-3.5 and GPT-4 perform at levels matching or exceeding those of children in structured ToM tasks, suggesting the spontaneous emergence of ToM-like behavior as models scale in size. Through mechanistic interpretation, Jamali et al. [2023] offer evidence of parallels between the LLM embeddings and neurons in the human brain.
However, these claims have been challenged. Ullman [2023], Shapira et al. [2023a], Holterman and Deemter [2023] show that minor task changes often cause LLMs’ performance to drop, suggesting reliance on brittle heuristics rather than true semantic understanding. To address this, researchers have developed more robust benchmarks and evaluation protocols featuring procedurally generated, higher-order, or broader ToM tasks [Gandhi et al., 2023, Chen et al., 2024d, He et al., 2023, Xu et al., 2024a, Shinoda et al., 2025]. These benchmarks frequently reveal significant declines in performance on complex tasks, such as 6th-order belief attribution [Street et al., 2024] or faux pas detection [Strachan et al., 2024a].
Some argue these failures stem from a lack of general commonsense reasoning rather than an inability to represent mental states [Pi et al., 2024]. Riemer et al. [2025] further suggest distinguishing between literal and functional ToM, noting that current benchmarks inadequately assess functional ToM. While newer models continue to improve on structured ToM tasks [Strachan et al., 2024a, Street et al., 2024], they remain inconsistent in open-ended, adversarial, or pragmatic reasoning scenarios [Sclar et al., 2024, Yu et al., 2025, Nickel et al., 2024]. Comparisons with human baselines show that LLMs can approximate ToM behavior in narrow contexts but still fall short of general human-like social cognition [Strachan et al., 2024a, van Duijn et al., 2023, Jones et al., 2024]. Recent work has also extended ToM evaluation to multimodal settings [Chen et al., 2024c, Strachan et al., 2024b, Jin et al., 2024a] and multi-agent interactions [Li et al., 2023b].
Main findings in ToM.
@@
EI is the subset of social intelligence that involves the ability to monitor one’s own and others’ feelings and emotions, to discriminate among them and to use this information to guide one’s thinking and actions [Salovey and Mayer, 1990].
Recent work introduces benchmarks to assess EI in LLMs using structured, theory-driven tasks. Tools like EmoBench [Sabour et al., 2024] and EQ-Bench [Paech, 2023] operationalize constructs such as emotional understanding and regulation, often referencing frameworks like Mayer-Salovey-Caruso Emotional Intelligence Test. Other studies adapt established psychometric instruments, including Situational Evaluation of Complex Emotional Understanding (SECEU) [Wang et al., 2023b], the Levels of Emotional Awareness Scale (LEAS) [Elyoseph et al., 2023], the Toronto Alexithymia Scale (TAS-20) and Empathy Quotient (EQ-60) [Patel and Fan, 2023], among others [Huang et al., 2024b]. Advanced models like GPT-4 often match or surpass human baselines in emotional awareness and understanding [Wang et al., 2023b, Elyoseph et al., 2023, Patel and Fan, 2023], though they lack deep reflexive analysis of emotional experience and motivation [Vzorinab et al., 2024]. Recent benchmarks also extend EI evaluation to multimodal settings [Hu et al., 2025].
Main findings in EI.